• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Annand

    Log entry time 08:01:53 on December 2,2008

    Entry number 251714

    keyword=Owl Shift Summary

    0.00 Continue production running 12uA on 3He,
    Field 25 Gauss transverse.
    LHRS @ 16 deg., 2.35 GeV/c, polarity -ve
    Beam lamda/2 OUT


    0.00 Run 4104 Continuation from swing (started 23.46) Q = 0.0392 C

    Happex 30494 Continuation from swing

    0.40 Run 4105 Production Q = 0.0392 C

    1.55 Run 4106 Production Q = 0.0389 C

    2.47 Beam continues to be unstable >15 trips last 1/2 hour
    However the position on the BPMs is stable
    A(0.15, 1.6)
    B(-1.5, 1.2)

    3.05 Run 4107 Very short...ended as forgot to change Happex run
    Negligible Q

    3.05 Happex run 30495 started

    3.06 Run 4108 Production Q = 0.0392 C

    4.10 Run 4109 Production Q = 0.0393 C

    5.12 Run 4110 Production Q = 0.0389 C

    6.21 Run 4111 Production Q = 0.0392 C

    7.26 Run 4112 Production Q = 0.0047 C

    7.36 Beam stopped for start of MCC studies

    7.41 Happex run 30495 ended

    Accumulated production-beam charge 0.279 C